### v1

@@ B.2 Parameter comparison
Here we explore the effects that alternative parameters have on the outcome of the RMHD simulations. Figure 7 showcases the influence of the magnetic field configuration on the EVPAs and total intensity, for a fixed . In the left panel the poloidal magnetic field configuration produces an EVPA pattern that is at a constant angle to the bulk jet flow, therefore disagreeing with the observations. Furthermore, the right panel illustrates the effects of a helical magnetic field configuration on the jet flow. Here the EVPAs exhibit some rotation, but again it does not match the observed one. The Stokes contours exhibit some bifurcation but only downstream, not reaching all the way into the compact core. In summary, the toroidal magnetic field configuration (see Fig. 1) reproduces the observed linearly polarised structure more faithfully than the other two configurations that we investigated.
As a next step, we then constrained the magnetic field configuration to be toroidal and alter . In Fig. 8 we present the effect of increasing on the simulated linearly polarised signatures. The left panel corresponds to and the right one to . We find that in this case the EVPA patterns retain their bimodality. The increase in also resulted in an apparent elongation of the limb brightened structure of the core, due to Doppler boosting, overall exceeding the extent of the sub-mas region of our observations. Therefore, both our RMHD simulations, as well as the observational evidence, are suggestive of only mildly relativistic velocities present in the sub-mas region of 3C 84.

### v2

@@ B.2 Parameter comparison
Here we explore the effects that alternative parameters have on the outcome of the RMHD simulations. Figure 7 showcases the influence of the magnetic field configuration on the EVPAs and total intensity, for a fixed . In the left panel the poloidal magnetic field configuration produces an EVPA pattern that is at a constant angle to the bulk jet flow, and therefore disagreeing with the observations. Furthermore, the right panel illustrates the effects of a helical magnetic field configuration on the jet flow. Here the EVPAs exhibit some rotation, but again it does not match the observed pattern. The Stokes contours exhibit some bifurcation, but only downstream and not reaching all the way into the compact core. In summary, the toroidal magnetic field configuration (see Fig. 1) reproduces the observed linearly polarised structure more faithfully than the other two configurations that we investigated.
As a next step, we then constrained the magnetic field configuration to be toroidal and altered . In Fig. 8 we present the effect of increasing on the simulated linearly polarised signatures. The left panel corresponds to and the right panel to . We find that in this case the EVPA patterns retain their bimodality. The increase in also resulted in an apparent elongation of the limb-brightened structure of the core, due to Doppler boosting, overall exceeding the extent of the sub-mas region of our observations. Therefore, our RMHD simulations and the observational evidence are both suggestive of only mildly relativistic velocities present in the sub-mas region of 3C 84.